OpenStudy (richyw):

This might be stupid, but how do I factor \(z^4+1\). I can do it by calculating the roots the long way and then write, it out, but what are the algebraic steps that get me from \(z^4+1\) to\((z+i)(z-i)(z+1)(z-1)\)

OpenStudy (kinggeorge):

First off, those aren't the actual roots of \(z^4+1\). Second, you could notice that \(z^4+1\) is the difference of two squares.\[z^4+1=z^4-(-1)=(z^2-i)(z^2+i)\]Each of these factors is then also a difference of two squares.

OpenStudy (richyw):

right, I just realized that haha.

OpenStudy (usukidoll):

why are there 4 King George pictures?

OpenStudy (anonymous):

#dontquestionitgowithit

OpenStudy (usukidoll):

King Friday.. Daniel Striped Tiger..Mr. Rogers Neighborhood

OpenStudy (shadowlegendx):

I love KingGeorge To make this post relevant I will say that math is awesome and I have no idea how to solve this question but good luck ;)

OpenStudy (compassionate):

You can put it in binomials then separate it into a trinomial.

OpenStudy (kinggeorge):

The second step is then\[(z^2-i)(z^2+i)=(z+\sqrt{i})(z-\sqrt{i})(z+\sqrt{-i})(z-\sqrt{-i}).\]If you were to write out \(\sqrt{i}\) and \(\sqrt{-i}\) in their exponential forms, you would get a possibly nicer looking factorization.

OpenStudy (primeralph):

There are other ways, but that's a quick method.

OpenStudy (richyw):

yes thank you! I was looking for the quick method. The way I did it was to write \[z^4=e^{i(\pi/2+2\pi n)}\]and work from there, but I was just curious. Thank you.

OpenStudy (richyw):

oops!

OpenStudy (richyw):

\(e^{i( \pi+2\pi n)}\)

OpenStudy (kinggeorge):

That's another way you could do it. Then\[z=e^{i(\pi+2\pi n)/4}=e^{i\pi/4+i\pi n/2}.\]Every time \(n\) increases by 4, you get the same value. So the roots will be \[e^{i\pi/4+i\pi/2},e^{i\pi/4+i\pi},e^{i\pi/4+i\pi 3/2},e^{i\pi/4}\]

OpenStudy (kinggeorge):

Put more simply, those are just\[e^{i\pi/4},e^{3i\pi/4},e^{5i\pi/4},e^{7i\pi/4}.\]